OpenStudy (anonymous):

From a base elevation of 7100 ft, a moutain peak in Colorado rises to summit elevation elevation of 14,622 ft over a horizontal distance of 15,848 ft. Find the grade of the peak.

OpenStudy (saifoo.khan):

G = Grade = Slope.

OpenStudy (saifoo.khan):

\[G = \frac{elevation}{\text{horizontal distance}}\]

OpenStudy (saifoo.khan):

\[ \frac{14622 - 7100}{15848}\to \frac{7522}{15848} \to \frac{3761}{7924} \to 0.475\]

OpenStudy (anonymous):

Is that 0.475%, because the question ask me for the percent? Thanks

OpenStudy (saifoo.khan):

nope, then i would be something like, 47.46% idk how much we have to round it off.

OpenStudy (anonymous):

it says round to the nearest tenth as needed

OpenStudy (saifoo.khan):

47.5%

OpenStudy (anonymous):

ok thanks

OpenStudy (saifoo.khan):

np

OpenStudy (anonymous):

I don't know how many questions I am aloud to ask but, can you answer this? Find the slope of this line (-3,1) (2,-3) m=?

OpenStudy (saifoo.khan):

@shadowfiend , why can't i tag this user in comments box?

OpenStudy (saifoo.khan):

@babygirl, you are allowed to ask as many questions as u want. :)

OpenStudy (saifoo.khan):

Slope = y2- y1 / x2- x1 Slope = -4/5